OpenStudy (anonymous):

Find inverse function for: f(x)=In(9x+5) Steps would be great

OpenStudy (anonymous):

\[\begin{array}{rcc}\text{Let y = f(x)} \\ \implies y = \ln(9x+5) \\ \implies e^{y}= 9x +5 \\ \implies \frac{e^{y}-5}{9}= x\\ \text{Hence,} & f^{-1}x= \frac{e^{x}-5}{9}\end{array}\]

OpenStudy (anonymous):

Thanks, Upsilon. Very comprehensive work shown to get the solution. I appreciate you time.

OpenStudy (anonymous):

your*